OpenStudy (anonymous):

the positive square root of 80 is ?

OpenStudy (anonymous):

4 sqrt 5

OpenStudy (accessdenied):

\[ \begin{split} \sqrt{80} &= \sqrt{4 * 20}\\ &= \sqrt{4 * 4 * 5}\\ &= \sqrt{4} *\sqrt{4} *\sqrt{5}\\ &= 2 * 2 * \sqrt{5}\\ &= 4 \sqrt{5}\\ \end{split} \]

OpenStudy (anonymous):

i would strongly advise - if not asked to - do not change to decimal.